A 0.010-kg bullet is fired from a 5.0-kg gun with a muzzle velocity of 230 m/s. Call the direction in which the bullet is fired from the +x direction .

Part A:
While the bullet is traveling in the barrel,what is the ratio of the acceleration of the gun to the acceleration of the bullet ?

Part B:
What is the speed of the bullet relative to the ground at the instant the bullet leaves the barrel?
1
Expert's answer
2020-04-20T19:06:28-0400

"m_1v_1=m_2v_2\\0.01*230=5*v_2"

"v_2=0.46 m\/s"


As the momentum on both sides is same and lets assume that the momentum change is in "\Delta t" time

then force is equal to "\frac{\Delta P}{\Delta t }=ma"

"m_1a_1=m_2a_2"

"\frac{m_1}{m_2}=\frac{a_2}{a_1}"

"\frac{a_2}{a_1}=\frac{5}{0.01}=500"